Area is 27 of a triangle
Height is 3x the base
what is the base

Answer by Maths68(1474)

We know
Area of triangle = (base*height)/2
Given:
Area=a=27
Base=b=?
Hieght=h=3b
Area of triangle = (base*height)/2
27=(b*3b)/2
27*2=b*3b
54=3b^2
Or
3b^2=54
b^2=54/3
b^2=18
b= √18
b=4.24
Base=4.24